1) Suppose a shell is shot from a cannon inclined at an angle of 45° with the ground and the shell is given an initial velocity of 1000 ft/sec.

a) What are the horizontal and vertical velocities of the shell?

b) Write the parametric equations of the motion of the shell?

c) Find the range of the shell.

d) What is the maximum height above ground attained by the shell during its flight?
Answer #1

a)

Horizontal velocity, Vx = 1000 cos 45 = 707.1 m/s

vertical velocity, Vy = 1000 sin 45 = 707.1 m/s

=====

b)

y = x tan 45 - 0.5 g x^2 / ( v^2 (cos 45)^2)

y = x - 0.0322 x^2

======

c)

R = u^2 sin (2 x) / g = 1000^2 / 32.2 = 3.1*10^4 ft

======

d)

max height

H = Vy^2 / 2g = 7763.8 ft

======

Comment in case any doubt, will reply for sure . Goodluck